# Morphology-Specific Peptide Discovery via Masked Conditional Generative Modeling

### v2

@@ Methods
$\mathrm{AP}\;=\;\frac{\bar{S}_{\mathrm{last2}}}{\bar{S}_{\mathrm{first2}}}.$ (11)
For RMOI, we identify, for the final saved frame of the trajectory, the largest peptide aggregate using a periodic-boundary, surface-to-surface connectivity graph. Each coarse-grained bead $j$ is assigned a radius $r_{j}$ and a mass $m_{j}$ (by MARTINI bead class, using defaults [45]). Two beads $j,k$ are considered connected if their minimum-image center distance $d_{jk}$ satisfies
@@
[…]
*Acknowledgements Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was funded by the ERC (StG SupraModel) - 101077842.
[…]
Author contributions Costa, N.: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Visualization. Zavadlav, J.: Writing - Review & Editing, Supervision.
[…]
*Competing interests The authors declare no competing interests.

### v3

@@ 4 Methods
$\mathrm{AP}\;=\;\frac{\bar{S}_{\mathrm{first2}}}{\bar{S}_{\mathrm{last2}}}.$ (12)
For RMOI, we identify, for the final saved frame of the trajectory, the largest peptide aggregate using a periodic-boundary, surface-to-surface connectivity graph. Each coarse-grained bead $j$ is assigned a radius $r_{j}$ and a mass $m_{j}$ (by MARTINI bead class, using defaults souzaMartini3General2021). Two beads $j,k$ are considered connected if their minimum-image center distance $d_{jk}$ satisfies
@@
Statistical Analysis
For screening-yield comparisons (hits under a finite sampling budget), we report binomial proportions with two-sided Wilson score 95% confidence intervals. Screening efficiency is summarized by calls per hit defined as $N/H$, where $N$ is the number of generated sequences and $H$ is the number of proxy-compliant hits (reported as $\infty$ when $H=0$).
To compare PepMorph candidates against aggregation-matched descriptor-fail control cohorts, we use two-sided Fisher exact tests on $2\times 2$ contingency tables (success/failure by cohort) and report the corresponding odds ratios. When a zero cell occurs, odds ratios are interpreted qualitatively; exact $p$-values from Fisher’s test are reported.
{acknowledgement}
We thank the early access to the SuperMUC-NG Phase 2 cluster, allowing us to do extensive control CG-MD simulations, which considerably strengthened our work.
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was funded by the ERC (StG SupraModel) - 101077842.
@@
Author contributions
Costa, N.: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Resources, Data Curation, Writing - Original Draft, Visualization. Zavadlav, J.: Writing - Review & Editing, Supervision.
{suppinfo}
The following files are available free of charge.
  • supplementary_material.pdf: Supplementary material, containing further dataset analysis and composition plots; all hyperparameters of the models used; analysis of the generation, including sampling costs and efficiency curves; per-peptide CG-MD visual snapshots and corresponding metrics and assessment of all runs; description and analysis of the robustness cohorts, including CG-MD visual snapshots and corresponding metrics and assessment of all runs; description and results of the executed atomistic simulations.